What rate gives ₹ 280 as interest on a sum of ₹ 56,000 in 2 years?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given: -P = ₹56000, SI = ₹280, t = 2 years.
We know that,
R = (100 × SI) / (P × T)
= (100 × 280)/ (56000 × 2)
= (1 × 28) / (56 × 2)
= (1 × 14) / (56 × 1)
= (1 × 1) / (4 × 1)
= (1/ 4)
= 0.25%

Thanks